OpenStudy (anonymous):

-2x+10y=-22 -2x-4y=6

OpenStudy (marissalovescats):

Elimination

OpenStudy (marissalovescats):

What can you multiply the bottom equation by to cancel out the x's?

OpenStudy (anonymous):

u munis it

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

multiply the bottom equation by -1 to cancel out the xs

OpenStudy (anonymous):

itll give you -2x + 10y = -22 2x +4y = -6 14y = -28 y = -2

OpenStudy (anonymous):

-2x -4y = 6 -2x -4(-2) = 6 -2x +8 = 6 -2x = 6-8 -2x = -2 x = 1

OpenStudy (anonymous):

whoops shoulda just made this a whole answer but test it by -2(1) + 10(-2) = -22 and it works woot

OpenStudy (anonymous):

awesome face

OpenStudy (marissalovescats):

Yup he's got it lol. Sorry I was helping someone else, otherwise I woulda answered but good job! :)